OpenStudy (anonymous):

3x + 5/16= 3/4-1/8x-1/2 Every time I work this out I get -5. But the back of my book says its -1/50 what am I doing wrong?

OpenStudy (amistre64):

(3x + 5/16= 3/4-1/8x-1/2) * 16 48x + 5= 12 - 2x -8 46x = -9 you will need to better define the setup using grouping symbols or the equation editor.

OpenStudy (amistre64):

and was spose to be 50x :)

OpenStudy (amistre64):

ugh, in any event, i forgot how to add and subtract; but it looks like the book is defining the setup differently than you have presented

OpenStudy (amistre64):

ah, got it .... you get -5, the book gets -1/50 and i get -1/50 when i add and subtract it correctly

OpenStudy (amistre64):

without knowing the steps you took, its hard to say what might have gone wrong

OpenStudy (amistre64):

\[3x + \frac 5{16}= \frac 34-\frac 18x-\frac 12\] since 2,4,8,16 are all common factors of 16, lets clear the fractions by multiplying this by 16 \[16\left(3x + \frac 5{16}= \frac 34-\frac 18x-\frac 12\right)\] \[3(16)x + \frac {5(16)}{16}= \frac {3(16)}4-\frac {1(16)}8x-\frac {1(16)}2\] \[48x + 5= 12-2x-8\]

OpenStudy (amistre64):

make sure you simplified the fractions correctly after multiplying by 16

OpenStudy (amistre64):

after that, its just adding 2x, subtracting 5, and dividing off 50 to get to x

OpenStudy (anonymous):

Oh, I forgot to multiply the 16 to the 3. thank you!

OpenStudy (amistre64):

keep up the good work ;)